# State-action control barrier functions: Imposing safety on learning-based control with low online computational costs

### v1

@@ V-D Discussion on state-action CBFs
One limitation of the parameterization (7) is that it sacrifices the universal approximation property [28] of NNs, since it restricts $Q_{\theta}$ to be quadratic on $u$. This may result in a large error bound $\Delta$ and cause constraint violations according to Theorem 2. As our main motivation for introducing state-action CBFs is to improve online computational efficiency, such a sacrifice is acceptable. Conversely, fully parameterizing $Q_{\theta}$ with a continuous NN, possessing universal approximation capabilities, increases the online computational complexity and introduces challenges in finding the global optimum of (15). In conclusion, when applying an approximation to the state-action CBF, there is typically a natural trade-off between online computational efficiency and safety.
Our proposed concept of state-action CBFs is inspired by the definition of state-action (Q) value functions in reinforcement learning [29]. Both kinds of functions do not only characterize the energy of each state but also quantify the quality of taking each action in each state. This similarity indicates that it is possible to use reinforcement learning methods [30] such as Q value iteration and Q learning to learn state-action CBFs directly, circumventing the collection of CBF samples. This methodology is plausible as the optimal control problem (V-A) can be approached via dynamic programming [30]. However, we do not adopt RL methods to learn state-action CBFs because the approximation error could accumulate when the iteration progress goes on. To interpret this, as $\alpha\geq 1$ in (V-A), $\alpha$ leads to a non-contractive dynamic programming model and consequently, convergence of RL algorithms is in general not guaranteed.
@@ VI Case study
Safety rate CPU time Safety rate CPU time Safety rate CPU time
No filter 63.52% 0.024 ms 36.66% 0.024 ms 56.99% 0.023 ms
Standard CBF 63.52% 2.3 ms 39.93% 2.4 ms 57.17% 2.3 ms
Quadratic state-action CBF 70.05% 1 ms 47.19% 1.1 ms 65.52% 0.9 ms
Nonlinear state-action CBF 65.15% 2.5 ms 40.19% 3.0 ms 60.80% 2.9 ms
Quadratic tightened state-action CBF 98.73% 1.0 ms 98.73% 1.0 ms 92.92% 1.0 ms
Nonlinear tightened state-action CBF 82.40% 2.7 ms 82.03% 2.7 ms 90.56% 3.0 ms
Quadratic contractive state-action CBF 100% 0.9 ms 100% 1.0 ms 100% 0.9 ms
Nonlinear contractive state-action CBF 96.91% 2.1 ms 92.20% 2.0 ms 94.57% 3.0 ms
@@
We validate the proposed methods for an active inverted pendulum interacting with elastic walls, which is a piecewise affine system [3]. The detailed model can be found in Appendix -E. The simulations are conducted in MATLAB R2021a on an AMD Core R7-5800H CPU @3.20GHz machine. All optimization problems involved in the policy filters (II-B) and (15) are solved using the Matlab function “fmincon”. The problem (V-A) and the MPC problem are transformed equivalently into mixed-integer quadratic programming problems and are solved by Gurobi [31].
@@
The system contains the state $x=[\theta\;\dot{\theta}]^{T}$ where $\theta$ is the pendulum angle, and the input $u$, which is the torque acting on the bottom end of the pendulum. The state and input constraints are given by $|x_{1}|\leq 0.15$, $|x_{2}|\leq 1$, and $|u|\leq 4$. The system is discretized using the explicit Euler method with time step 0.05s.
@@
Using the approach in Appendix -D and letting $\lambda=0.2$ for Option 2 and $\lambda=0.05$ for Option 3, we obtain the initial CBF $B_{0}(x)=x^{T}Px-1$ with $P=\left[\begin{array}[]{rr}124.74&7.44\\ 7.44&2.24\end{array}\right]$, $\left[\begin{array}[]{rr}283.40&20.25\\ 20.25&4.69\end{array}\right]$, and $\left[\begin{array}[]{rr}522.39&25.69\\ 25.69&5.39\end{array}\right]$ for Options 1, 2, and 3, respectively. We sample from the state-input space $\left\{(x,u)\big{|}\;|x_{1}|\leq 0.16,|x_{2}|\leq 1.1,|u|\leq 4\right\}$ on a uniform grid and we obtain $40^{3}$ samples $\{(x_{i}^{\mathrm{s}},u_{i}^{\mathrm{s}})\}^{40^{3}}_{i=1}$. The successor state $f\left(x_{i}^{\mathrm{s}},u_{i}^{\mathrm{s}}\right)$ of each sample $\left(x_{i}^{s},u_{i}^{s}\right)$ is fed to the CBF generator (V-A) with $k=7$. According to Section V-B, only those samples satisfying $B_{k}(f(x_{i}^{\mathrm{s}},u_{i}^{\mathrm{s}}))\leq 10$ are obtained. After this procedure, 42109, 38609 and 37027 samples are collected in each option.
@@
We compare state-action CBFs and standard CBFs. We consider 4 different kinds of safety certificates: (i) State-action CBF for the original state constraint; (ii) State-action CBF for the tightened state constraint $|x_{1}|\leq 0.14$, $|x_{2}|\leq 0.9$; (iii) Contractive state-action CBF; (iv) Standard CBF for the original state constraint. For the learning, we use NNs with three hidden hyperbolic tangent layers containing 16, 64 and 8 neurons. We parameterize each standard CBF by an NN. For the parameterization of each state-action CBF, we use (7) or fully parameterize it by an NN. Besides, we also consider 4 different control policies. The first three include (i) supervised learning the MPC policy [19], (ii) the approximate dynamic programming (ADP) policy [3], which is one kind of RL policy, and (iii) the LQR policy for the linearized system around the origin. These policies may be unsafe and are thus taken as $\pi_{0}$ in (II-B) and (15). The learning-based control policies (i) and (ii) are also represented by NNs. All NNs are trained via the stochastic gradient descent algorithm [32]. The fourth one is implicit MPC. One can refer to Appendix -E for a detailed description of these policies.
@@
To test all the control methods, we randomly sample initial states from the state sample set $\{x_{i}^{\mathrm{s}}\}^{40^{3}}_{i=1}$ and select a total of 551 initial states $x$ that satisfy $B_{k}(x)\in(-0.1,0)$. These initial states are contained in the safe set of $B_{k}$, and thus can be steered to the origin without constraint violation. Meanwhile, since the values of $B_{k}$ taken at these states are close to zero, these initial states are near the boundary of the safe set. Therefore, they are in general more difficult to stabilize than any other states in the safe set. For each initial state, we simulate the closed-loop systems for 50 time steps.
The simulation results are shown in Table 1. The performance metrics include the rate of safety and the average CPU time for computing the control input per time step. The safety rate is defined as the number of initial states leading to a safe closed-loop trajectory divided by the total number (551) of initial states. Table 1 shows that directly applying the basic policies yields a lower rate of constraint satisfaction compared to passing these policies through safety filters. Among all types of CBFs, contractive state-action CBFs perform the best in the rate of constraint satisfaction, resulting in almost zero constraint violations. The superior performance of the quadratic contractive state-action CBF over its nonlinear counterpart with respect to constraint satisfaction is attributed to the emergence of suboptimal solutions to (15) when employing the nonlinear contractive state-action CBF. State-action CBFs with state constraint tightening make a significant improvement in safety performance compared to those CBFs without constraint tightening. Regarding the CPU time, we find that the implicit MPC controllers need the most time to compute the control inputs. Standard CBFs and state-action CBFs with a full NN parameterization have a comparable online computation time (from 2 to 3 ms per time step), while state-action CBFs with the parameterization (7) achieve a dramatic decrease to near 1 ms per time step. It should be noticed that the computational advantage of the quadratic state-action CBF will be larger if multi-start optimization is used to solve (II-B) or (15) with a fully parameterized CBF.
@@ VII Conclusions and future work
This paper has presented a new type of CBFs, called state-action CBFs, which can synthesize safe control policies for general constrained nonlinear systems with a very small online computational overhead. The paper developed a new approach to learn the proposed CBF from data. We also proposed a constraint tightening approach to improve the robustness of the learned CBFs to learning errors. Simulation on an inverted pendulum with hybrid dynamics shows that the proposed method can achieve zero constraint violations and can compute the control input in about 1 ms.
In the future, we will first focus on learning the proposed CBFs for large-scale problems in a computationally efficient manner. We will also investigate how the proposed CBFs can be exploited to guide the learning of the basic policy. This will help to reduce the potential degradation of other control performance measure caused by the safety filter. More importantly, as the proposed safety filter does not need any information about the model, it can be applied in model-free cases provided that the state-action CBFs can be learned without knowing the model. As model-free (reinforcement) learning control has received much attention recently, the state-action CBF approach is an important future research direction to achieve safe learning-based control in a completely model-free manner.

### v2

@@ V-D Discussion on SACBFs
One limitation of the parameterization (7) is that it sacrifices the universal approximation property [37] of NNs, since it restricts $Q_{\theta}$ to be quadratic on $u$. This may result in a large error bound $\Delta$ and cause constraint violations according to Theorem 2. As our main motivation for introducing state-action CBFs is to improve online computational efficiency, such a sacrifice is acceptable. Conversely, fully parameterizing $Q_{\theta}$ with a continuous NN, possessing universal approximation capabilities, increases the online computational complexity and introduces challenges in finding the global optimum of (V-C). In conclusion, when applying an approximation to the SACBF, there is typically a natural trade-off between online computational efficiency and safety.
Note that a parametrization of an SACBF $Q$, instead of parameterizing a standard CBF $B$, may increase the approximation error on $B(f)$, especially when $Q$ is parameterized in a quadratic form. This error can result in the learned $Q$ failing to satisfy the conditions of a valid SACBF. To address this issue, we have proposed the constraint tightening approach (Section IV) to ensure that the learning-based approach can still generate a valid SACBF under mild approximation errors. However, a quadratic parametrization of $Q$ typically requires more stringent constraint tightening, which can lead to increased conservatism in the learned barrier certificates.
The conservatism of the learned SACBF $Q_{\theta}$ as well as the policy $\pi_{\theta}$ is strongly influenced by $\Delta$. According to Theorem 2, to guarantee the recursive feasibility of (V-C) and the safety of $\pi_{\theta}$, a larger $\Delta$ requires a larger $\lambda$ in (V-A). This, in turn, increases the conservatism of the target $Q$. In practice, to minimize the conservatism, one can start with a small $\lambda$ and iteratively increase it until Assumption 1 and $\Delta\leq\lambda/2$ are satisfied. Increasing the approximation capability of the chosen function approximator, as discussed in [38], can also be beneficial.
We focus on discrete-time nonlinear systems because our work targets safety in learning-based control methods, like RL and supervised learning of MPC policies, which typically use discrete-time models [10, 25]. Additionally, safety filters for discrete-time systems are computationally more challenging, leading to nonlinear programs, whereas continuous-time systems involve simpler convex quadratic programs [22]. Therefore, the computational benefits of introducing SACBFs are more evident in the context of discrete-time nonlinear systems.
Compared to learning standard CBFs [17], the sampling complexity of learning the proposed SACBFs increases mildly because the domain of the function to be approximated only changes from $\mathbb{R}^{n_{x}}$ to $\mathbb{R}^{n_{x}+n_{u}}$. More importantly, advanced sampling strategies, such as active learning [36] and counterexample-guided inductive synthesis [39], can be utilized to improve the scalability of approximating $Q$.
@@ VI Case study
Safety rate CPU time Cost Safety rate CPU time Cost Safety rate CPU time Cost
No filter 63.52% 0.024 ms 99.69 36.66% 0.024 ms 125.11 56.99% 0.023 ms 122.10
Standard CBF 63.52% 2.3 ms 126.18 39.93% 2.4 ms 161.82 57.17% 2.3 ms 120.00
Quad. SACBF 70.05% 1 ms 131.76 47.19% 1.1 ms 161.83 65.52% 0.9 ms 115.80
Nonlin. SACBF 65.15% 2.5 ms 132.22 40.19% 3.0 ms 168.19 60.80% 2.9 ms 137.30
Quad tightened SACBF 98.73% 1.0 ms 97.29 98.73% 1.0 ms 99.44 92.92% 1.0 ms 91.47
Nonlin. tightened SACBF 82.40% 2.7 ms 117.26 82.03% 2.7 ms 116.56 90.56% 3.0 ms 116.92
Quad. contr. SACBF 100% 0.9 ms 98.84 100% 1.0 ms 101.91 100% 0.9 ms 93.10
Nonlin. contr. SACBF 96.91% 2.1 ms 110.44 92.20% 2.0 ms 128.99 94.57% 3.0 ms 109.97
@@
We validate the proposed methods for an active inverted pendulum interacting with elastic walls, which is a piecewise affine system [3]. The detailed model can be found in Appendix -E. The simulations are conducted in MATLAB R2021a on an AMD Core R7-5800H CPU @3.20GHz machine. All optimization problems involved in the policy filters (5) and (V-C) are solved using the Matlab function “fmincon” and the active-set optimization algorithm. The problem (V-A) and the MPC problem are transformed equivalently into mixed-integer quadratic programming problems and are solved using Gurobi [40].
@@
The system contains the state $x=[\theta\;\dot{\theta}]^{T}$ where $\theta$ is the pendulum angle, and the input $u$, which is the torque acting on the bottom end of the pendulum. The state and input constraints [^3] The theoretical analysis in the previous sections is applicable to nonlinear constraints. However, for this case study, we focus on linear constraints for illustrative purposes. are given by $|x_{1}|\leq 0.15$, $|x_{2}|\leq 1$, and $|u|\leq 4$. The system is discretized using the explicit Euler method with time step 0.05s.
@@
Using the approach in Appendix -D and letting $\lambda=0.2$ for Option 2 and $\lambda=0.05$ for Option 3, we obtain the initial CBF $B_{0}(x)=x^{T}Px-1$ with $P=\left[\begin{array}[]{rr}124.74&7.44\\ 7.44&2.24\end{array}\right]$, $\left[\begin{array}[]{rr}283.40&20.25\\ 20.25&4.69\end{array}\right]$, and $\left[\begin{array}[]{rr}522.39&25.69\\ 25.69&5.39\end{array}\right]$ for Options 1, 2, and 3, respectively. We sample from the state-input space $\left\{(x,u)\big{|}\;|x_{1}|\leq 0.16,|x_{2}|\leq 1.1,|u|\leq 4\right\}$ on a uniform grid and we obtain $40^{3}$ samples $\{(x_{i}^{(\mathrm{s})},u_{i}^{(\mathrm{s})})\}^{40^{3}}_{i=1}$. The successor state $f\left(x_{i}^{(\mathrm{s})},u_{i}^{(\mathrm{s})}\right)$ of each sample $\left(x_{i}^{(\mathrm{s})},u_{i}^{(\mathrm{s})}\right)$ is fed to the CBF generator (V-A) with $k=7$. Only those samples satisfying $B_{k}(f(x_{i}^{(\mathrm{s})},u_{i}^{(\mathrm{s})}))\leq 10$ are collected. After this procedure, 42109, 38609 and 37027 samples are collected in each option.
@@
We compare SACBFs and standard CBFs. We consider 4 different kinds of safety certificates: (i) SACBF for the original state constraint; (ii) SACBF for the tightened state constraint $|x_{1}|\leq 0.14$, $|x_{2}|\leq 0.9$; (iii) Contractive-set SACBF; (iv) Standard CBF for the original state constraint. For the learning, we use NNs with three hidden hyperbolic tangent layers containing 16, 64 and 8 neurons. We parameterize each standard CBF by an NN. For the parameterization of each SACBF, we use (7) or fully parameterize it by an NN. Besides, we also consider 4 different control policies. The first three include (i) supervised learning of the MPC policy [14], (ii) the approximate dynamic programming (ADP) policy [3], which is a kind of RL policy, and (iii) the LQR policy for the linearized system around the origin. These policies may be unsafe and are thus taken as $\pi_{0}$ in (5) and (V-C). The learning-based control policies (i) and (ii) are also represented by NNs. All NNs are trained via the stochastic gradient descent algorithm [33]. The main reason for choosing these three particular policies is that they can achieve sub-optimal control objectives using very limited computational resources. However, they generally cannot guarantee constraint satisfaction. Therefore, we use these three policies as the baseline to demonstrate that the proposed safety filter can enhance the rate of constraint satisfaction for these policies. The fourth control policy is implicit MPC. One can refer to Appendix -E for a detailed description of these policies.
@@
To test all the control methods, we uniformly randomly sample initial states from the state sample set $\{x_{i}^{(\mathrm{s})}\}^{40^{3}}_{i=1}$ and select a total of 551 initial states $x$ that satisfy $B_{k}(x)\in(-0.1,0)$. These initial states are contained in the safe set of $B_{k}$, and thus can be steered to the origin without constraint violation. Meanwhile, since the values of $B_{k}$ taken at these states are close to zero, these initial states are near the boundary of the safe set. Therefore, they are in general more difficult to stabilize than any other states in the safe set. For each initial state, we simulate the closed-loop systems for 50 time steps.
The simulation results are shown in Table 1. The performance metrics include the rate of safety, the average CPU time for computing the control input per time step, and the total cost $\sum_{t=0}^{50}C(x_{t},u_{t})$ with the stage cost the same as that in the considered basic controllers. The safety rate is defined as the number of initial states leading to a safe closed-loop trajectory divided by the total number of initial states. Overall, the table indicates that incorporating appropriate safety filters, especially quadratic and nonlinear SACBFs, can significantly improve the safety performance of control policies without excessively compromising on CPU time and total cost. The online computational benefit of using the SACBF with the quadratic parametrization (7) primarily stems from the fact that solving a convex quadratic program is typically faster than solving a nonlinear program with the same number of constraints and decision variables, especially when the nonlinear constraints involve an NN. The maximum performance loss is about 34%, which occurs when applying the nonlinear SACBF to the ADP controller. Considering that the main objective of our work is to impose safety using low computational resources, such a performance degradation is acceptable. The inferior performance of the nonlinear contractive-set SACBF compared to its quadratic counterpart with respect to constraint satisfaction is primarily due to the suboptimal solutions when solving (V-C). In other words, the assumption of perfect nonlinear optimization does not hold in practical scenarios, leading to the rare instances of constraint violations observed in Table 1. It should be noticed that the computational advantage of the quadratic SACBF will be larger if multi-start nonlinear optimization is used to solve (5) or (V-C) with a fully parameterized CBF.
(a) State trajectories and safe set for the SACBF.
(b) State trajectories and safe set for the contractive-set SACBF.
Figure 1: Closed-loop trajectories of the learning MPC controller under the safety filter (V-C).
In Figure 1, we demonstrate the state-trajectories of the closed-loop system with the policy $\pi_{\theta}$ from (V-C), with $\pi_{0}$ being the learning MPC controller and $Q_{\theta}$ being the nonlinear SACBF (left) or the nonlinear contractive-set SACBF (right). Without constraint tightening, the system evolution records several constraint violations, as evidenced by the 6th line of Table 1. Conversely, the contractiveness property of SACBFs ensures that almost all state trajectories remain within its safe set.
@@ VII Conclusions and future work
This paper has presented a new type of CBFs, called SACBFs, that can synthesize safe controllers for general constrained nonlinear systems with a very small online computational overhead. We have also developed a new approach to learn the proposed CBF from data, and proposed a constraint tightening approach to improve the robustness of the learned CBFs to learning errors.
In the future, we will first focus on learning the proposed CBFs for large-scale problems in a computationally efficient manner. We will also investigate how the proposed CBFs can be exploited to guide the learning of the basic policy. This will help to reduce the potential degradation of other control performance measure caused by the safety filter. We will investigate whether undesired equilibria might arise when applying the proposed SACBF-based safety filter, as observed for continuous-time systems. More importantly, as the proposed safety filter does not need any information about the model, it can be applied in model-free cases provided that the SACBFs can be learned without knowing the model. As model-free (reinforcement) learning control has received much attention recently, the SACBF approach is an important future research direction to achieve safe learning-based control in a completely model-free manner.